# Pixel-level Scene Understanding in One Token: Visual States Need What-is-Where Composition

### 14 Mar 2026 v1

@@ Abstract
For robotic agents operating in dynamic environments, learning visual state representations from streaming video observations is essential for sequential decision making. Recent self-supervised learning methods have shown strong transferability across vision tasks, but they do not explicitly address what a good visual state should encode. We argue that effective visual states must capture what-is-where by jointly encoding the semantic identities of scene elements and their spatial locations, enabling reliable detection of subtle dynamics across observations. To this end, we propose CroBo, a visual state representation learning framework based on a global-to-local reconstruction objective. Given a reference observation compressed into a compact bottleneck token, CroBo learns to reconstruct heavily masked patches in a local target crop from sparse visible cues, using the global bottleneck token as context. This learning objective encourages the bottleneck token to encode a fine-grained representation of scene-wide semantic entities, including their identities, spatial locations, and configurations. As a result, the learned visual states reveal how scene elements move and interact over time, supporting sequential decision making. We evaluate CroBo on diverse vision-based robot policy learning benchmarks, where it achieves state-of-the-art performance. Reconstruction analyses and perceptual straightness experiments further show that the learned representations preserve pixel-level scene composition and encode what-moves-where across observations.
@@ 1 Introduction
We argue that a visual state representation for sequential decision making must capture what-is-where by encoding both the semantic identities of scene elements and their precise spatial locations. For example, to recognize that a robot hand has moved from left to right across a sequence of observations, the representation must jointly encode the identity of the hand and its position in the scene; more broadly, it must encode which semantic contents exist in the observation and where they are located, so that even subtle spatial changes can be directly detected. From this perspective, understanding scene dynamics in robotics can be viewed as a form of pixel-level video understanding, where the representation must preserve spatial semantics while remaining sensitive to how they evolve across observations. We posit that this what-is-where understanding is the foundation for reasoning about temporal dynamics.
@@ 2 Method
Claim. Our goal is to learn a compact visual state representation that enables dynamics-aware scene understanding and is transferable to real-world video applications (e.g., robot learning). To this end, we posit that such a representation must capture what-is-where scene composition: it should preserve which semantic entities are present in the scene, where they are located, how they relate semantically to each other, and how they are arranged within the overall scene composition. This enables reliable detection of subtle changes and interactions across observations—i.e., understanding what-moves-where in the scene.
@@ 3.1 Implementation details
Pre-training. For a fair comparison, we follow the pre-training setup of RSP [22]. Specifically, we use ViT-S/16 [9] and pre-train on the Kinetics-400 dataset [24] for 400 epochs with a repeated sampling value of 2 [20, 12]. For spatial views $\mathrm{x}^{g}$ and $\mathrm{x}^{l}$, we adopt the global-to-local cropping configuration of [11]. We apply a masking ratio of 90% to the target view $\mathrm{x}^{l}$. The decoder consists of 8 layers with an embedding dimension of 512. We optimize the model using AdamW [26] with a batch size of 1536.
Competitors. We compare our method with (1) standard self-supervised learning (SSL) methods, including MAE [17], DINO [5], and DINOv2 [28]. We also compare with (2) SSL methods for dynamic scenes, including SiamMAE [15], CropMAE [11], RSP [22], and ToBo [25].
@@ 3.2 Vision-based Robot Learning
In this section, we evaluate the proposed method on two vision-based policy learning benchmarks, Franka Kitchen and the DeepMind Control Suite (DMC), covering robotic manipulation and locomotion tasks in simulated environments.
@@
Table 2: Scalability of our method on Franka Kitchen [14]. Success rates (%) are reported for ViT-S/16, ViT-B/16, and ViT-L/16 backbones pre-trained on Kinetics-400 [24] for 100 epochs. While ViT-S results are provided for CroBo, ViT-B and ViT-L scales include comparisons with SiamMAE [15], RSP [22], and ToBo [25]. $\Delta$ denotes the performance margin relative to the best-performing baseline in each scale. Best results are in bold and second best are underlined. CroBo results are averaged over 10 trials, while baseline results and those marked with are from [25].
@@
Scaling behavior. To evaluate the scalability of our method across different model capacities, we benchmark CroBo against baselines using ViT-B/16 and ViT-L/16 backbones, and report ViT-S/16 results for CroBo. All models are pre-trained on Kinetics-400 [24] for 100 epochs and evaluated on Franka Kitchen [14]. As shown in Tab. 2, CroBo consistently outperforms SiamMAE [15], RSP [22], and ToBo [25] across all architecture scales. The base and large variants achieve 5-task average success rates of 70.5% and 71.1%, respectively, surpassing the prior state of the art by substantial margins of +9.4% and +7.8%. Remarkably, even our smallest backbone (ViT-S/16) attains a 65.0% average success rate, outperforming all baselines that use the much larger ViT-L/16 architecture. These results demonstrate that the performance gains of CroBo are not due to model scale. Instead, they reflect a fundamentally stronger representation that generalizes robustly across the full spectrum of model capacities.
Figure 4: Reconstruction of CroBo. We visualize image reconstructions on CLEVR [23] (columns 1-3), DAVIS [29] (columns 4-5), and Franka Kitchen [14] (column 6). Using the bottleneck token of the source view as context, CroBo reconstructs a highly masked (90%) target view cropped from the source view. The results show that CroBo representations effectively capture the global scene structure, including object identities, locations, and their spatial relationships.
@@ 3.3 Qualitative Analysis
In this section, we examine how well the bottleneck token captures what-is-where scene composition through reconstruction visualizations.
Target datasets. We visualize image reconstructions on CLEVR [23], DAVIS [29], and Franka Kitchen [14]. CLEVR provides synthetic scenes with simple spatial layouts and clearly defined object attributes (e.g., color, shape and materials), making it well-suited for evaluating what-is-where scene composition. DAVIS extends the analysis to natural video scenes with diverse semantic objects and appearance variations, whereas Franka Kitchen provides complex robotic manipulation scenes.
Analysis. We follow the same reconstruction setup used during training (Fig. 4). A target view is obtained by randomly cropping the reference image. We then mask 90% of the target patches and reconstruct the view from the visible patches using the bottleneck token as context. The first example shows that the bottleneck token preserves object attributes together with spatial locations and relationships: the two cyan spheres are correctly reconstructed even though they are completely invisible in the masked input. The second example shows that the representation captures fine-grained object and scene details, such as metallic material and shadows. The fourth example shows that the bottleneck token retains the person-horse relationship in the scene: from only a few visible person parts, the model plausibly reconstructs both the horse shape and person’s pose. This suggests that the bottleneck token captures how objects are semantically arranged in the scene. The last example helps explain the strong robotic performance of CroBo. The representation consistently preserves what-is-where scene composition even in complex scenes, which is critical for localizing affordance-bearing objects during manipulation. Additional results are provided in the supplementary material.
@@ 3.4 Perceptual Straightness in Video
In this section, we analyze the temporal geometry of representations of our dynamics-aware SSL model. Specifically, we investigate whether representations evolve smoothly over time when processing natural videos using the concept of perceptual straightening [18, 19]. In particular, we measure the local curvature of representation trajectories induced by sequential video frames and compare them across models.
Figure 5: Perceptual straightness of representation dynamics in video. (a) Local curvature of representation trajectories measured on DAVIS videos. Curvature is computed as the angle between consecutive representation differences, averaged across videos. Lower curvature indicates smoother temporal dynamics. CroBo consistently exhibits lower curvature than prior models such as DINOv2 and CropMAE, suggesting more temporally coherent representations that better preserve what-moves-where across frames. (b) Representation trajectories across video frames visualized using PCA. Each point corresponds to the representation of a frame, and colors indicate temporal progression. CroBo produces smooth and locally linear trajectories that follow the natural evolution of the scene, whereas DINOv2 and CropMAE produce irregular and tangled trajectories.
Perceptual straightness. Perceptual straightness is the property of perceptual representations in which the visual system transforms inputs to follow a straight path in the representation space. Since observations vary in a complex and highly nonlinear way in the pixel space, it is infeasible to make meaningful predictions from raw observations [27]. Therefore effective video understanding systems [18, 1] compress raw observations into a visual state that has a locally straight temporal trajectory. This makes predicting future states possible via linear extrapolation.
Measuring perceptual straightness. Local curvature is the main metric to assess the perceptual straightness of representations [18, 21]. Formally, let $z_{t}=f_{\theta}(x_{t})$ denote the representation of frame $x_{t}$ produced by an encoder $f_{\theta}$. The temporal trajectory of a video in representation space is given by the sequence $\{z_{t}\}_{t=1}^{T}$. We measure the local curvature using the angle between consecutive difference vectors $\Delta z_{t}=z_{t+1}-z_{t}$ and $\Delta z_{t+1}=z_{t+2}-z_{t+1}$ as
@@
Lower curvature values indicate that representation trajectories evolve more linearly over time, suggesting temporally consistent features that effectively encode what-moves-where. In other words, a model understanding smooth motion will produce a smooth trajectory of representations [18]. We evaluate the perceptual straightness of representations generated by models both quantitatively and qualitatively to demonstrate the effectiveness of our approach.
@@
Quantitative comparison. We measure local curvature on videos of the DAVIS validation set. For each video containing at least 50 frames, we compute the trajectory and the local curvature for the first 50 frames and plot the average local curvature at Fig. 5(a). CroBo achieves an average curvature of $75.4^{\circ}$, compared to $103.28^{\circ}$ of DINOv2. The lower curvature values signal that representations of our dynamics-aware SSL model follows a locally linear path. This result suggests that our model is able to capture subtle temporal differences between adjacent frames and maintain a coherent representation of what-moves-where.
@@
An example trajectory. As an intuitive example, we compute representations of frames from the kite-walk video in the DAVIS dataset [29]. Then we compute two principal components of the trajectory via PCA and plot it at Fig. 5(b) following [1]. While DINOv2 and CropMAE generated highly jagged trajectories, CroBo produced a smooth trajectory, demonstrating a consistent understanding of what-moves-where. The trajectory is also compliant to the contents of the video; while the person moves right and then left in the video, the representation also moves back and forth in the first principal component.
3.5 Ablation on Target View Construction
In this section, we compare our global-to-local relationship with another promising source-to-target relationship for visual state representation learning (SRL).
Preliminary. To learn visual correspondences between two views, SiamMAE [15] constructs a current-to-future relationship, by randomly sampling two frames from a video. In contrast, CropMAE [11] constructs a global-to-local relationship, by generating the target view through cropping the source view. In our ablation study, we denote these two formulations as Time and Crop, respectively. Our proposed method corresponds to Crop.
@@ 3.5 Ablation on Target View Construction
Time vs Crop. We compare the two source-to-target formulations, Time and Crop, by pre-training ViT-S/16 on Kinetics-400 for 100 epochs and evaluating on Franka Kitchen. As shown in Tab. 3, Crop consistently outperforms Time across all five tasks. This result is somewhat surprising, since Crop does not use temporal cues, yet achieves better transfer on dynamic downstream tasks that require temporal awareness. We attribute this to two factors. First, Crop provides a cleaner and more tractable supervisory signal: because the source view fully contains the target view, the reconstruction target is well-defined and spatially grounded. In contrast, Time relies on temporally shifted targets and therefore introduces additional ambiguity from object motion, camera motion, and appearance change across frames. Second, the scene structure learned from Crop serves as a stronger inductive bias for understanding dynamics. This is consistent with our hypothesis that explicitly learning what-is-where is crucial for capturing how the scene evolves over time (i.e., what-moves-where).
@@
Table 3: Ablation on the construction of the target view. Models are pre-trained on Kinetics-400 for 100 epochs and evaluated on Franka Kitchen (success rate %). “Time” uses a future frame at time $t+k$ as the target view. “Crop” uses a cropped local region as the target view. Our method is highlighted in gray.
@@
Time + Crop. Given that Time provides temporal cues while Crop offers clear spatial grounding, one may expect that combining the two would yield a stronger supervisory signal. However, Tab. 3 shows that Time+Crop performs worse than both individual formulations. We conjecture that this degradation stems from a conflict between the two source-to-target relationships. When combined, the model must simultaneously resolve spatial localization and temporal change, making the target more ambiguous and reducing the quality of the supervisory signal. This further supports our use of Crop-based global-to-local formulation.
@@ 4 Related Work
Self-supervised Learning (SSL) aims to learn rich and transferable visual representations from unlabeled images or videos. A large body of SSL methods build on representation alignment across augmented views, using contrastive objectives [31, 6, 16, 8, 4] or teacher-student self-distillation [5, 13, 33, 28] to learn invariant features. Meanwhile, reconstruction-based approaches learn representations by recovering masked pixels or tokens [2, 17, 33, 32]. Despite their success, most SSL methods have primarily been evaluated on traditional tasks such as image classification and segmentation, leaving transferable representation for dynamic scene understanding largely underexplored.
Siamese Masked Autoencoder [15], on the other hand, targets dynamic scene understanding by learning visual correspondence from videos. It employs a Siamese encoder [3] to predict masked patches in one frame from another, thereby learning patch-wise correspondence across time. CropMAE [11] further shows that similar correspondence can be learned even from static images by reconstructing one cropped view from another. However, their patch-wise correspondence learning is suboptimal for downstream video understanding tasks that require a single compact frame representation, such as robotic manipulation.
State Representation Learning (SRL), as introduced by ToBo [25], addresses the above limitation by learning to encode an observation into a single compact visual state (i.e., the bottleneck token). It learns this state by reconstructing heavily masked target patches from the bottleneck token of a temporally adjacent frame and a few visible target patches. This encourages temporal relationships to be preserved in the representation, but does not explicitly account for within-frame spatial structure and scene composition.
Our work extends this line of work by studying a dynamics-aware SSL method for learning state representations that capture scene composition. Like prior works, we use a Siamese network to process paired views and learn from their relationship through image reconstruction [15, 11, 25]. However, our work differs in two key aspects. First, we identify that compact visual state representations for dynamic world understanding should preserve scene composition. Second, to achieve this, we propose a novel global-to-local reconstruction framework that learns to encode scene composition into the bottleneck token.

### 25 Mar 2026 v2

@@ Abstract
For robotic agents operating in dynamic environments, learning visual state representations from streaming video observations is essential for sequential decision making. Recent self-supervised learning methods have shown strong transferability across vision tasks, but they do not explicitly address what a good visual state should encode. We argue that effective visual states must capture what-is-where by jointly encoding the semantic identities of scene elements and their spatial locations, enabling reliable detection of subtle dynamics across observations. To this end, we propose CroBo, a visual state representation learning framework based on a global-to-local reconstruction objective. Given a reference observation compressed into a compact bottleneck token, CroBo learns to reconstruct heavily masked patches in a local target crop from sparse visible cues, using the global bottleneck token as context. This learning objective encourages the bottleneck token to encode a fine-grained representation of scene-wide semantic entities, including their identities, spatial locations, and configurations. As a result, the learned visual states reveal how scene elements move and interact over time, supporting sequential decision making. We evaluate CroBo on diverse vision-based robot policy learning benchmarks, where it achieves state-of-the-art performance. Reconstruction analyses and perceptual straightness experiments further show that the learned representations preserve pixel-level scene composition and encode what-moves-where across observations. Project page available at: https://seokminlee-chris.github.io/CroBo-ProjectPage.
@@ 1 Introduction
We argue that a visual state representation for sequential decision making must capture the what-is-where scene composition by encoding both the semantic identities of scene elements and their precise spatial locations. Here, what-is-where denotes whether the representation retains which semantic entities are present in the scene and how they are spatially located and arranged within the overall scene composition. For example, as illustrated in Fig. 1, to detect that a robot hand moved from right to left across observations, the representation must encode both the identity of the hand and its position, so that even subtle spatial changes can be directly detected. From this perspective, understanding scene dynamics in robotics can be viewed as a form of pixel-level video understanding, where the representation must preserve spatial semantics while remaining sensitive to how they evolve across observations.
@@ 2 Method
Claim. Our goal is to learn a compact visual state representation that enables dynamics-aware scene understanding and is transferable to real-world video applications (e.g., robot learning). To this end, we posit that such a representation must capture what-is-where scene composition: it should preserve which semantic entities are present in the scene, and how they are semantically arranged within the overall scene composition. This enables reliable detection of subtle changes and interactions across observations—i.e., understanding what-moves-where in the scene.
@@ 3.1 Implementation details
Pre-training. For a fair comparison, we follow the pre-training setup of RSP [37]. We use ViT-S/16 [17] and pre-train on the Kinetics-400 dataset [41] for 400 epochs with a repeated sampling value of 2 [34, 20]. For spatial views $\mathrm{x}^{g}$ and $\mathrm{x}^{l}$, we adopt the global-to-local cropping configuration of [19], with global and local crop scales of [0.5, 1.0] and [0.3, 0.6], respectively. We apply a masking ratio of 90% to the target view $\mathrm{x}^{l}$. The decoder consists of 8 layers with an embedding dimension of 512. We optimize the model using AdamW [45] with a batch size of 1536. See Appendix. A for full implementation details.
Competitors. We compare our method with standard SSL methods (MAE [31], DINO [7], and DINOv2 [47]) and dynamic-scene SSL methods (SiamMAE [24], CropMAE [19], RSP [37], and ToBo [42]).
@@ 3.2 Vision-based Robot Learning
In this section, we evaluate our method on two vision-based policy learning benchmarks, Franka Kitchen and the DeepMind Control Suite (DMC), covering robotic manipulation and locomotion tasks in simulated environments.
@@
Table 2: Scalability of our method. Success rates (%) are reported for ViT-S/16, ViT-B/16, and ViT-L/16 backbones pre-trained on Kinetics-400 [41] for 100 epochs. While ViT-S results are provided for CroBo, ViT-B and ViT-L scales include comparisons with SiamMAE [24], RSP [37], and ToBo [42]. $\Delta$ denotes the performance margin relative to the best-performing baseline in each scale. Best results are in bold and second best are underlined. CroBo results are averaged over 10 trials, while baseline results and those marked with are from [42].
@@
Scaling behavior. To evaluate scalability across model capacities, we benchmark CroBo against baselines using ViT-B/16 and ViT-L/16 backbones. All models are pre-trained on Kinetics-400 [41] for 100 epochs and evaluated on Franka Kitchen [23]. As shown in Tab. 2, CroBo consistently outperforms SiamMAE [24], RSP [37], and ToBo [42] across all architecture scales. The base and large variants achieve 5-task average success rates of 70.5% and 71.1%, respectively, surpassing the prior state of the art by substantial margins of +9.4% and +7.8%. Remarkably, even our smallest backbone (ViT-S/16) achieves a 65.0% average success rate, outperforming all baselines built on the much larger ViT-L/16. This shows that the gains of CroBo arise not from model scale, but from a stronger representation that generalizes across architectures of varying capacity.
@@ 3.3 Qualitative Analysis
In this section, we examine how well the bottleneck token captures what-is-where scene composition through reconstruction visualizations. We follow the same reconstruction setup used during training (Fig. 4).
Target datasets. We visualize image reconstructions on CLEVR [39], DAVIS [49], MOSEv2 [16], and Franka Kitchen [23]. CLEVR provides synthetic scenes with simple layouts and well-defined object attributes (e.g., color, shape, and materials), enabling evaluation of what-is-where scene composition. DAVIS considers natural video scenes with a dominant object, whereas MOSEv2 features more complex and crowded dynamic scenes with multiple interacting instances. Franka Kitchen provides complex robotic manipulation scenes.
Figure 4: Reconstruction of CroBo. We visualize image reconstructions on CLEVR [39], DAVIS [49], MOSEv2 [16] and Franka Kitchen [23]. Using the bottleneck token of the reference view as context, CroBo reconstructs a highly masked (90%) target view cropped from the reference view. The results show that the bottleneck token from the reference view provides sufficient information to recover overall scene structure, including object identity, location, and spatial relationships. (e.g., in column 1, the two cyan spheres are not visible in the masked input, yet their positions are correctly reconstructed using the reference bottleneck token).
Analysis. The first column shows that the bottleneck token preserves object attributes along with spatial locations: the two cyan spheres are accurately reconstructed, even though they are fully occluded in the masked input. The second column further demonstrates that the representation retains fine-grained object and scene details, including metallic reflectance and shadows. The third through sixth columns show that these properties remain robust in more complex and cluttered scenes. In the third column, CroBo successfully recovers the horse’s overall shape and the rider’s pose from only a few visible human parts. Moreover, the fourth through sixth columns illustrate that the model faithfully recovers multiple scene elements and their complex configurations. These results suggest that the bottleneck token encodes a coherent scene representation with a strong understanding of semantic arrangement even under increased complexity. This capability contributes to the strong robotic performance of CroBo by capturing reliable what-is-where scene composition during manipulation. Additional results are provided in Appendix C.
@@ 3.4 Perceptual Straightness in Video
In this section, we analyze the temporal geometry of representations of our dynamics-aware SSL model. Specifically, we investigate whether representations evolve smoothly over time when processing natural videos using the concept of perceptual straightening [32, 33].
Figure 5: Perceptual straightness of representation dynamics in video. (a) Local curvature of representation trajectories measured on DAVIS videos. Curvature is computed as the angle between consecutive representation differences, averaged across videos. Lower curvature indicates smoother temporal dynamics. CroBo consistently exhibits lower curvature than prior models such as DINOv2 and CropMAE, suggesting more temporally coherent representations that better preserve what-moves-where across frames. (b) Representation trajectories across video frames visualized using PCA. Each point corresponds to the representation of a frame, and colors indicate temporal progression. CroBo produces smooth and locally linear trajectories that follow the natural evolution of the scene compared to other models.
Perceptual straightness. Perceptual straightness is the property of perceptual representations in which the visual system transforms inputs to follow a straight path in the representation space. Since observations vary in a highly nonlinear way in the pixel space, it is infeasible to make predictions from raw observations [46]. Therefore effective video understanding systems [32, 1] compress raw observations into a visual state that has a locally straight temporal trajectory, enabling future state prediction via extrapolation.
Measuring perceptual straightness. Local curvature is the main metric to assess the perceptual straightness of representations [32, 36]. Let $z_{t}$ denote the representation of frame $x_{t}$. The temporal trajectory of a video in representation space is the sequence $\{z_{t}\}_{t=1}^{T}$. We measure the local curvature using the angle between consecutive difference vectors $\Delta z_{t}=z_{t+1}-z_{t}$ and $\Delta z_{t+1}=z_{t+2}-z_{t+1}$ as
@@
Lower curvature indicates that representation trajectories evolve more linearly over time, suggesting temporally consistent features that effectively encode what-moves-where. In other words, a model understanding smooth motion will produce a smooth trajectory of representations [32].
@@
Quantitative comparison. We measure local curvature on videos of the DAVIS validation set and compare across models. For each video containing at least 50 frames, we compute the trajectory and the local curvature for the first 50 frames and plot the average local curvature at Fig. 5(a). CroBo achieves an average curvature of $75.4^{\circ}$, compared to $103.28^{\circ}$ of DINOv2. The lower curvature signal that representations of our model follow a locally linear path. This suggests that CroBo is able to capture subtle temporal differences between adjacent frames and maintain a coherent representation of what-moves-where.
@@
An example trajectory. As an intuitive example, we evaluate representations of frames from the kite-walk video in DAVIS [49], compute two principal components of the trajectory via PCA, and plot it at Fig. 5(b) following [1]. While DINOv2 and CropMAE generated highly jagged trajectories, CroBo produced a smooth trajectory, compliant to the contents of the video. While the person moves right and then left in the video, the representation also moves back and forth in the first principal component. Additional results are provided in Appendix. D.
3.5 Ablation Studies
3.5.1 Ablation on Target View Construction
In this section, we compare our global-to-local formulation with alternative source-to-target relationships for visual state representation learning (SRL). Following SiamMAE [24], we consider a current-to-future relationship in which two frames are sampled from a video; we refer to this formulation as Time. In contrast, our method adopts a global-to-local relationship [19], where the target view is obtained by randomly cropping the source view; we refer to this formulation as Crop. We also consider a hybrid variant, Time + Crop, in which the target is constructed by cropping a temporally subsequent frame.
@@ 3.5.1 Ablation on Target View Construction
Time vs. Crop. We compare Time and Crop by pre-training ViT-S/16 on Kinetics-400 for 100 epochs and evaluating on Franka Kitchen. As shown in Tab. 3(a), Crop consistently outperforms Time across all five tasks. We attribute this result to the quality of the supervisory signal. Under Crop, the target is spatially grounded because it is fully contained in the source view, making reconstruction well-defined. By contrast, Time introduces ambiguity from object motion, camera motion, and appearance changes across frames. These results suggest that learning spatial scene composition through Crop provides a stronger basis for visual state representation than relying on temporal correspondence.
Time + Crop. One may expect that combining temporal variation with spatial grounding would provide a richer learning signal. However, Tab. 3(a) shows that Time+Crop performs worse than both Time and Crop. We conjecture that combining the two makes the prediction target less coherent: the model must jointly account for spatial localization and temporal change, which increases target ambiguity and weakens supervision. This result further supports our choice of the Crop-based global-to-local formulation.
3.5.2 Ablation on Masking ratio
We vary the masking ratio of the target scene from 75% to 95% to verify whether limiting target-view information encourages the decoder to more effectively exploit the visual scene representation encoded from the source scene. As shown in Sec. 3.5.2, performance improves substantially as the masking ratio increases from 75% to 90%, supporting our hypothesis that reducing target-side clues encourages the model to make fuller use of the compressed reference information. Moreover, although our default setting is 90%, a more aggressive masking ratio of 95% yields the best performance across all tasks. These results indicate that further limiting the visible target content encourages the model to draw more effectively on the source scene, leading to the best overall performance.
@@ 3.5.2 Ablation on Masking ratio
Table 3: Ablation studies on target view design and masking ratio. Models are pre-trained on Kinetics-400 for 100 epochs and evaluated on Franka Kitchen (success rate %). Our method is highlighted in gray.
@@
(a) Target view construction
Masking Ratio Knob1 Light Sdoor Ldoor Micro
75% 41.4 70.6 94.0 22.6 35.0
[…]
95% 59.0 86.6 99.4 41.2 58.0
(b) Masking ratio
@@ 4 Related Work
Self-supervised Learning (SSL) aims to learn rich and transferable visual representations from unlabeled images or videos. A large body of SSL methods build on representation alignment across augmented views, using contrastive objectives [54, 9, 30, 11, 6] or teacher-student self-distillation [7, 22, 61, 47] to learn invariant features. Meanwhile, reconstruction-based approaches learn representations by recovering masked pixels or tokens [3, 31, 61, 57]. Despite their success on traditional tasks such as image classification and segmentation, transferable representations for dynamic scene understanding remain underexplored.
Siamese Masked Autoencoder [24], on the other hand, targets dynamic scene understanding by learning visual correspondence from videos with a Siamese encoder [5]. It predicts masked patches in one frame from another, learning patch-wise correspondence across time. CropMAE [19] shows that similar correspondence can be learned even from static images by reconstructing one cropped view from another. However, such patch-wise correspondence is suboptimal for downstream video tasks that require a single compact frame representation, such as robotic manipulation.
State Representation Learning (SRL), introduced by ToBo [42], addresses the above limitation by learning to encode an observation into a single compact visual state (i.e., the bottleneck token). It learns this state by reconstructing heavily masked target patches from the bottleneck token of a temporally adjacent frame and a few visible target patches. This encourages temporal relationships to be preserved in the representation, but does not explicitly account for within-frame spatial structure and scene composition.
Our work extends this line of work by studying a dynamics-aware SSL method for learning state representations that capture scene composition. While we follow prior works in using a Siamese network and paired-view reconstruction [24, 19, 42], our work differs in two key aspects. First, we explicitly argue that compact visual states for dynamic world understanding should preserve scene composition. Second, to achieve this, we propose a novel global-to-local reconstruction framework that learns to encode scene composition into the bottleneck token.